OpenStudy (anonymous):

Please help! Kurt says that you can take the square root of each side of an equation. Therefore, he says that (x+2)^2 + (9y-5)^2+36 and (x+2) + (y-5)=6 are equivalent equations. Dana says the equations are not equivalent. Who is right? Explain your answer.

OpenStudy (anonymous):

you mean (x+2)^2 + (9y-5)^2= 36 or (x+2)^2 + (9y-5)^2+36

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Dreaaaaaa

OpenStudy (anonymous):

I didn't even notice I put the 9! I meant \[(x+2)^{2}+(y-5)^{2}=36 and (x+2)+(y-5)=6\]

OpenStudy (anonymous):

well,first let's work with the second one... \(\color{lime}{ (x+2)+(y-5)=6 \Longrightarrow ((x+2) + y-5)^2 = 36} \)

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

can u find \( \large \color{blue}{((x+2) + y-5)^2} \)

OpenStudy (anonymous):

i don't know how to start off :/

OpenStudy (anonymous):

and sry i missed a bracket : \( \large \color{blue}{((x+2) + (y-5))^2} \)

OpenStudy (anonymous):

ok let's just see what would be \( (a+b)^2 \) you can find it in this way: \( (a+b)(a+b) \)

OpenStudy (anonymous):

it would be \( a^2 + b^2 + 2(a+b) \)

OpenStudy (anonymous):

so would it be \[x ^{2}+y ^{2}+7\]? Sorry i'm honestly so bad at math haha

OpenStudy (anonymous):

no no ... imagine that (x+2) = a and (y-5) = b and it would be : \(\color{red}{ (x+2)^2 + (y-5)^2 + 2(x+2)(y-5)} \)

OpenStudy (anonymous):

so ,\( \color{red}{ (x+2)^2 + (y-5)^2 + 2(x+2)(y-5) = 36} \) but we had \(\large \color{blue}{(x+2)^{2}+(y-5)^{2}=36} \)

OpenStudy (anonymous):

so they're not equivalent

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes! sorry i was trying to write it down and see if i got it haha thank you!